OpenStudy (anonymous):

Simplify 4 sqrt sign 500 over 4 sqrt sign 5

OpenStudy (anonymous):

is the four the index, as in \(\sqrt[4]{500}\) or is it a coefficient as in \(4\sqrt{500}\) ?

OpenStudy (anonymous):

The first one. The index.

OpenStudy (anonymous):

hmm then you can do this \[\sqrt[4]{500}=\sqrt[4]{5\times 100}=\sqrt[4]{5}\sqrt[4]{100}=\sqrt[4]{5}\sqrt[4]{10^2}=\sqrt[4]{5}\sqrt{10}\]

OpenStudy (anonymous):

or actually it would be a lot easier to write \[\frac{\sqrt[4]{500}}{\sqrt[4]{5}}=\sqrt[4]{100}=\sqrt{10}\]

OpenStudy (anonymous):

You made it seem easy xD. I appreciate your help!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

uh oh, I was just looking at the multiple choices here. and it's:

OpenStudy (anonymous):

How should I go about it now?

OpenStudy (anonymous):

can you post a shot of the question?

OpenStudy (anonymous):

Yeah, one second.

OpenStudy (anonymous):

hmm looks like that is a 400 up top, and not a 500

OpenStudy (anonymous):

That would probably help,wouldnt it?

OpenStudy (anonymous):

now the first step is to divide \(400\) by \(5\) and get \(\sqrt[4]{80}\)

OpenStudy (anonymous):

since \(80=16\times 5\) and \(\sqrt[4]{16}=2\) you are left with \[2\sqrt[4]{5}\]

OpenStudy (anonymous):

Once again. Thanks so much for your help. Haha.

OpenStudy (anonymous):

yw (again)